OpenStudy (anonymous):

What is the sum of a 12-term arithmetic sequence where the last term is 13 and the common difference is -10? 605 660 748 816

OpenStudy (campbell_st):

well you need to find the 1st term \[T_{n} = a + (n - 1)d\] is the formula to find a term in an arithmetic series You know the 12th term = 13, the common difference is -10 and the number of terms n = 12 so \[13 = a + (12 -1) \times(-10)\] solve for a. then the sum of the terms can be found using \[S_{n} = \frac{n}{2}[ a + l]\] n = number of terms, a = 1st term and l = last term hope this helps.